# Nevanlinna Theory of Algebroid Functions on Complete Kähler Manifolds

### v10

@@ 4.3.2. Calculus Lemma and Logarithmic Derivative Lemma
which gives for So, we conclude that
@@
holds for where is the Riemannian area element of
@@
holds for outside where is defined by
@@
Differentiating to obtain
In further
Applying Borel’s lemma to the left hand side on the above equality twice, then we see that there exists a set of finite Lebesgue measure such that
holds for outside Meanwhile, Corollary 4.16 means that there exists a constant such that
@@
holds for outside where is defined by
@@
holds for outside where is defined by
@@ 5.1. First Main Theorem
In what follows, we introduce Nevanlinna’s functions of algbroid functions. We are given a -valued algebroid function with branching set on which is defined by an irreducible algebraic equation
such that defines a meromorphic mapping with indeterminacy locus As we note that are the distinct single-valued meromorphic components of on outside Assume that
The characteristic function of for is defined by
@@
are called the proximity function, counting function of for respectively. For define
@@
According to the Vieta’s theorem, we obtain Combining Dynkin formula, it yields that
@@
In particular, and Similarly, we obtain
which leads to
@@
5.2. Estimate of Branching Divisor
Let be a -valued algebroid function on We use the same notations as defined in Section 5.1. Our purpose is to obtain an estimate of the branching divisor of For each we use to denote the all -valued, -valued components of at respectively. As seen before, we have Namely, has branching order at with and has branching order at The branching divisor of associated to is defined to be the divisor on satisfying that
The counting function of branching divisor of for is defined by
@@ 5.2. Estimate of Branching Divisor
For any we use to denote the all distinct function elements of at with leaves respectively. It is noted that From the assumption, we deduce that Since the set of all singular branching points of has complex codimension not smaller than it is sufficient to treat the case where is a non-singular branching point of By Theorem 2.14, we may write for
in a local holomorphic coordinate near such that has a local holomorphic coordinate If then is not a pole of Through a simple observation of we deduce from (29) immediately that there exist terms that contain the factor i.e., terms that contain the factor in For the worst case that for all we deduce that the factor appears a total of at least times in Hence,
@@
Theorem 5.3 (Estimate of branching divisor).
@@ 5.3. Second Main Theorem
Let be the -leaf complex manifold by a -valued algebroid function on which is a (non-compact) complex manifold of complex dimension due to Theorem 2.20. Let be the natural projection. Let be the pullback of complete Kähler metric on Note that is a complete degenerate Kähler metric on because could be not positive-definite at points in the branching set of Applying Corollary 3.8, can lift to a meromorphic function such that Note that is an analytic set of complex dimension Next, we explain that Dynkin formula still holds under the pullback metric
Set Picking a sequence of neighborhoods of such that
Let be a partition of the unity subordinate to the open covering of Then, and Fix a complete Kähler metric on Let Evidently, is a complete Kähler metric on with on Since then one may assume that (otherwise, a subsequence can be chosen so that as ). Let be the Laplace-Beltrami operators of on respectively. We have weakly as Take an arbitrary bounded domain with smooth boundary in Fix a point with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, and again by the harmonic measures of on with respect to respectively. It is not difficult to show that as and (weakly) uniformally on as due to
as for a Borel subset of Similarly, we have (weakly) uniformally on as in which stand for the volume measures of respectively. Set and Using Dynkin formula, we are led to
for a -class function with on outside a polar set of singularities at most. Hence, it yields by letting that
This means that Dynkin formula holds under the metric Furthermore, it is not hard to check that Theorem 4.12 and Theorem 4.20 also hold for under the corresponding curvature conditions, since is an analytic set of complex codimension
5.3.1. with non-negative Ricci curvature
Assume that is non-parabolic. With the pullback metric is also non-parabolic, with non-negative Ricci curvature. Let be the Kähler form of Take where is a bounded domain defined in Section 4.2.1. Recall the Green function for and the harmonic measure on defined in Section 4.2.1. Then, the characteristic function of is defined by
@@
Assume that We denote by the connected components of centered at respectively, with for Again, define
Then, it is easy to infer that is the Green function of for with a pole at satisfying Dirichlet boundary condition, as well as that is the harmonic measure on with respect to Set
@@
Note that is a -sheeted analytic covering, and two distinct connected components () may cross only at those points in the branching set of If is a -order branching point of then there are multi-leaf function elements with subordinate to equation (7). As we know that Since a point in may be a branching point of it yields from the definition of that
@@
Combining the above with (5.3.1), it follows that for any there exists a subset of finite Lebesgue measure such that
@@
5.3.2. with non-positive sectional curvature
Assume that has non-positive sectional curvature satisfying (3). Note that the curvature of also satisfies (3) under the pullback metric In the similar way, we obtain notations and by setting Applying the same arguments as in the proof of Theorem 5.4, we can show that
@@
It is known that is a complete non-compact Kähler manifold of constant sectional curvature under this metric. From (31), the associated Laplace-Beltrami operator is
@@
Hence, we have the minimal positive global Green function of for with a pole at 0:
where denotes the standard Euclidean area of the unit sphere in By definition (see Section 2.1), we have
@@
with Let be the Riemannian area element of By the second formula in (33) and volume form formula, it gives immediately
@@
Substituting (32) into the expression of and computing the integral, it is not hard to deduce that
We proceed to estimate with By
@@ 6.1. Propagation of Algebraic Dependence
If is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2), then denote by
@@
Applying Dynkin formula, we deduce that
This completes the proof. ∎
@@
Let be a -valued algebroid function in It suffices to show that on Set
Let be the diagonal of Since on then Let be the natural projection on the -th factor with Consider the holomorphic line bundle over where is the point line bundle over If on then there is a holomorphic section of over such that and Taking then one can have and On the other hand, implies that
@@ 6.2. Five-Value Type Theorem
Assume that is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2).
@@
which implies holds. ∎
@@
Assume that has non-positive sectional curvature satisfying (3).
@@
Assume that is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2).
@@
in the sense of currents, where and are understood as divisors. Thus, we obtain
@@
But, it contradicts with the assumption for This completes the proof. ∎
@@
Assume that has non-positive sectional curvature satisfying (3).

### v11

@@ 4.3.2. Calculus Lemma and Logarithmic Derivative Lemma
which implies that for all Thus, we conclude that
@@
holds for all where is the Riemannian area element of
@@
holds for all outside where is defined by
@@
Differentiating to get
In further, we have
Applying Borel’s growth lemma to the left hand side on this equality twice, we can deduce that there exists a set of finite Lebesgue measure such that
holds for all outside Meanwhile, Corollary 4.16 means that there exists a constant such that
@@
holds for all outside where is defined by
@@
holds for all outside where is defined by
@@ 5.1. First Main Theorem
In what follows, we introduce Nevanlinna’s functions of algbroid functions. Let be a -valued algebroid function with branch set on which is defined by an irreducible algebraic equation
such that defines a meromorphic mapping with indeterminacy locus In the following, we assume that
The characteristic function of on is defined by
@@
are called the proximity function, counting function of on respectively. For define
@@
According to Vieta’s theorem, we have
Since is polar, it yields from Dynkin formula that
@@
In particular, we have
Similarly, we obtain
which yields that
@@
5.2. Estimate of Branch Divisors
Let be a -valued algebroid function on We use the same notations as defined in Section 5.1. Our purpose is to obtain an estimate of the branch divisor of For each we use to stand for the all -valued, -valued components of at respectively. Then, we have and has branch order at with and has branch order at The branch divisor of associated to is defined by a divisor on such that
The counting function of branch divisor of for is defined by
@@ 5.2. Estimate of Branch Divisors
For any we use to denote the all distinct function elements of at with leaves respectively. It is noted that From the assumption, we deduce that Since the set of all singular branch points of has complex codimension not smaller than it is sufficient to deal with the case where is a non-singular branch point of By Theorem 2.14, we may write for
in a local holomorphic coordinate near such that has a local holomorphic coordinate If then is not a pole of Through a simple observation of we deduce from (29) immediately that there exist terms that contain the factor i.e., terms that contain the factor in For the worst case where for all we deduce that the factor appears a total of at least times in Hence, we get
@@
Theorem 5.3 (Estimate of branch divisors).
@@ 5.3. Second Main Theorem
Let be the -leaf complex manifold by a -valued algebroid function on which is of complex dimension by Theorem 2.20. Let be the natural projection. We equip with the pullback metric induced from via Note that is a positive semi-definite Kähler metric on i.e., is a Kähler metric on and degenerate on Since is a polar set with Dynkin formula still works on for the following reasons. Set Picking a sequence of neighborhoods of such that
Let be a partition of the unity subordinate to the open covering of Then, and Fix a Kähler metric on with Kähler form Set Since is positive definite on we see that is also positive definite on Moreover, since we conclude that is a Kähler metric on with on Let denote the Riemannian measures of respectively. By and as we can deduce that weakly as Let stand for the Laplace-Beltrami operators of on respectively. According to and weakly as one has weakly as Taking a bounded domain with smooth boundary in Fixing with Denote by the Green functions of for with a pole at satisfying Dirichlet boundary condition respectively, here is well-defined on Also, we denote by the harmonic measures of on with respect to respectively, here is well-defined on Since is a polar set, the previous arguments can show that in the sense of distributions and weakly as By Dynkin formula
for a -class function with on outside a polar set of singularities at most. Since is polar, we conclude by letting that
This means that Dynkin formula works on under the pullback metric By Corollary 2.21, one can lift to a meromorphic function such that In further, it is not difficult to examine that Theorem 4.2, Theorem 4.12 and Theorem 4.20 can apply to on each leaf of under the corresponding conditions.
5.3.1. has non-negative Ricci curvature
Assume that is non-parabolic. With the pullback metric one notes that is non-parabolic, with non-negative Ricci curvature outside Let be the Kähler form of Set where is the bounded domain defined in Section 4.2.1. We recall the definitions of Green function for and harmonic measure on in Section 4.2.1. Then, the characteristic function of is defined by
@@
Assume that Let be the connected components of centered at respectively, with that for each Let denote the Laplace-Beltrami operator of which is well-defined on Define
In a sense, we can regard outside that is the Green function of for with a pole at satisfying Dirichlet boundary condition, and that is the harmonic measure on with respect to Set
@@
Note that is a -sheeted analytic covering, and two distinct connected components of meet only at the points in Moreover, any point in is possibly a branch point of or equivalently, any point in is possibly a ramification point of Hence, it follows from that
@@
Combining the above with (5.3.1) and Theorem 5.3, it yields that for any there exists a subset of finite Lebesgue measure such that
@@
5.3.2. has non-positive sectional curvature
Assume that has curvature satisfying (3). Note that the curvature of also satisfies (3) outside under the pullback metric Similarly, we obtain notations and by setting Employing the similar arguments as in the proof of Theorem 5.4, we can conclude from Theorem 4.20 that
@@
It is known that is a complete non-compact Kähler manifold of constant sectional curvature under this metric. Using (31), the associated Laplace-Beltrami operator is
@@
Hence, we obtain the minimal positive global Green function of for with a pole at 0:
where stands for the standard Euclidean area of the unit sphere in By the definition (see Section 2.1), we have
@@
with Let be the Riemannian area element of By the second formula in (33), it gives immediately
@@
Substituting (32) into the expression of and computing the integral, it is not hard to deduce that We proceed to estimate with By
@@ 6.1. Propagation of Algebraic Dependence
If is non-parabolic with non-negative Ricci curvature and volume growth satisfying (2), then we denote by
@@
Applying Dynkin formula, we conclude that This completes the proof. ∎
@@
Let be a -valued algebroid function in It suffices to show that on Set Let denote the diagonal of Since on we obtain Let be the natural projection on the -th factor with We consider the holomorphic line bundle over where is the point line bundle over If on then there exists a holomorphic section of over such that and Take then one can have and On the other hand, implies that
@@ 6.2. Five-Value Type Theorem
Assume that is non-parabolic with volume growth satisfying (2).
@@
which implies that holds. ∎
@@
Assume that has curvature satisfying (3).
@@
Assume that is non-parabolic with volume growth satisfying (2).
@@
in the sense of currents, where and are understood as divisors. Thus, it yields from that
@@
which contradicts with the condition (6) for This completes the proof. ∎
@@
It is not hard to show that holds with under the condition for Employing Theorem 6.11, we have holds. Similarly, we have holds with Hence, we prove ∎
@@
Assume that has curvature satisfying (3).